How would you solve this problem? I need step by step on how it was solved please. 0.3x-0.2y=4
0.4+ 0.5y= -55over 7
What is the solution of the system?

Answer by stanbon(75887)

How would you solve this problem? I need step by step on how it was solved please. 0.3x-0.2y=4
0.4+ 0.5y= -55over 7
What is the solution of the system?
-----------------------------
Get rid of those pesky decimals.
-----------
3x - 2y = 40
4x + 5y = 79
------------
Multiply the 1st Eq. by 5 ; Multiply the 2nd Eq. by 2
---
15x - 10y = 200
8x + 10y = 158
-----------------------
23x = 358
x = 15.56..
---
Subtitute to solve for "y":
3(15.56) - 2y = 40
-2y = -6.68
y = 3.34
----------------------------------
Cheers,
Stan H.